Unknown gas x has relative density 1.517 according to air.What will be relative density of gas x according to helium? Mr(He)=4 g/mol,Mr(air)=29 g/mol
1
Expert's answer
2020-05-20T13:43:53-0400

Solution:

RD = Msubstance / Mreference,

where:

RD is relative density

Msubstance is the molar mass of the substance

Mreference is the molar mass of the reference


RD according to air = M(X) / M(air)

M(X) = RD according to air × M(air) = 1.517 × 29 g/mol = 43.993 g/mol

M(X) = 43.993 g/mol


RD according to helium = M(X) / M(He)

RD according to helium = (43.9993 g/mol) / (4 g/mol) = 10.998

RD according to helium = 10.998


Answer: Unknown gas X has relative density 10.998 according to helium (He).